I) In an equilibrium mixture at 100°C the following gas concentrations were found:



CO : 3.76 x 10-3 mol L-1 H2: 4.30 x 10-3 mol L-1





CH3OH: 4.17 x 10-10 mol L-1



Calculate the value of K(eq) at 100°C

1
Expert's answer
2022-02-14T17:05:07-0500

Solution:

CO(g) + 2H2(g) ⇌ CH3OH(g) + 52 kJ


The equilibrium constant (Keq) expression is as follows:



[CH3OH] = 4.17×10-10 M

[CO] = 3.76×10-3 M

[H2] = 4.30×10-3 M


Therefore,



Thus,

Keq = 6.00×10-3 M-2 = 6.00×10-3 L2 mol-2


Answer: The value of Keq is 6.00×10-3 L2 mol-2